a certain club has 15 male members. how many ways are there to form committees of 3
Answer by stanbon(75887)

a certain club has 15 male members. how many ways are there to form committees of 3
------
Ans: 15C3 = (15*14*13)/(1*2*3) = 5*7*13 = 35*13 = 350+90+15 = 455 ways
------------
Cheers,
Stan H.
-----------